How do you solve the following word problem?
Shirley is currently twice as old as HArry. Four years ago, Shirley was 3 times as old as Harry. How old were Shirley and Harry 4 years ago?
I can't remember how to solve the equation when there are two unknowns.

Answer by Paul(988)

Shirley's age = 2x
Hary's age = x
4 years ago Shirley's age was = 2x-4
Hary's age was = x-4
EQUATION:
3(x-4)=2x-4
3x-12=2x-4
x=8
2(8)=16
8-4=4
16-4=12
Hence, SHirley was 12 years old, and Hary was 4 years old four years ago.
Paul.